# A Survey of Self-Evolving Agents: On Path to Artificial Super Intelligence

### v3

@@ 6.2 Specialized Domain Evolution
The power of self-evolving agents extends directly to practical applications like coding, where their ability to autonomously adapt and improve offers a transformative approach to software development. SICA (robeyns2025self, ) demonstrates that a self-improving coding agent can autonomously edit its own codebase and improve its performance on benchmark tasks. EvoMAC (hu2024self, ) introduces a self-evolving paradigm on multi-agent collaboration networks, which automatically optimizes individual agent prompts and multi-agent workflows, significantly improving code generation performance by overcoming the limitations of manually designed systems. AgentCoder (huang2024agentcoder, ) also focuses on a multi-agent code generation framework that self-evolves through iterative refinement. A programmer agent continuously improves code based on feedback from a test executor agent, validated against independent test cases from a test designer, significantly boosting effectiveness and efficiency. Zhang et al. (zhang2025adaptive, ) enable LLM agents to continuously evolve by filtering high-quality answers, stratifying earned experiences by difficulty, and adaptively selecting demonstrations from self-generated data, leading to significant performance improvements and the construction of ML libraries. While these instances differ in their specific mechanisms—ranging from single-agent self-editing to complex multi-agent collaborative networks and experience-based learning—they commonly share the core principle of iterative self-improvement and autonomous adaptation to enhance coding capabilities. These advancements highlight how self-evolving agents can dramatically enhance coding efficiency and code quality by continuously learning and optimizing.
@@
Self-evolving GUI agents extend LLM capabilities from pure text reasoning to direct manipulation of desktop, web, and mobile interfaces, where they must cope with large discrete action spaces, heterogeneous layouts, and partial visual observability. Yuan et al. couple pixel-level vision with self-reinforcement, enabling the agent to iteratively refine click–type grounding accuracy without additional human labels (yuan2025enhancing, ). On real desktop software, the Navi agent from WindowsAgentArena replays and critiques its own failure trajectories, ultimately doubling its task-completion rate across 150 Windows challenges (bonatti2024windows, ). For open-web automation, WebVoyager fuses screenshot features with chain-of-thought reflection; successive self-fine-tuning raises its end-to-end success on unseen sites from 30 % to 59 % (he2024webvoyager, ), while ReAP adds episodic memories of past outcomes, recovering a further 29-percentage-point margin on previously failed queries (azam2025reap, ). Beyond RL and memory, AutoGUI continuously mines functionality annotations from live interfaces to expand a reusable skill library each training cycle (li2025autogui, ), and MobileUse deploys a hierarchical self-reflection stack that monitors, verifies, and revises smartphone actions in situ (wu2025mobileuse, ). Collectively, these systems epitomize the full triad of self-evolution— what evolves (grounding modules, skill memories), when it evolves (offline consolidation vs. online reflection), and how it evolves (reinforcement learning, synthetic data, hierarchical monitoring)—charting a path toward universally competent interface agents.
@@
The primary bottleneck in customizing agents for specialized domains like financial tasks lies in efficiently constructing and integrating a domain-specific knowledge base into the agent’s learning process—a challenge that can be effectively mitigated by incorporating self-evolving mechanisms. QuantAgent (wang2024quantagent, ) proposed a two-layer framework that iteratively refines the agent’s responses and automatically enhances its domain-specific knowledge base using feedback from simulated and real-world environments. This iterative process helps the agent progressively approximate optimal behavior, reduces reliance on costly human-curated datasets, and demonstrably improves its predictive accuracy and signal quality in trading tasks. TradingAgents (xiao2024tradingagents, ) incorporates dynamic processes such as reflection, reinforcement learning, and a feedback loop from real-world trading results, alongside collaborative debates, to continuously refine its strategies and enhance trading performance. These developments underscore the potential of self-evolving agents to revolutionize the financial domain by autonomously building domain expertise, adapting to dynamic market conditions, and continuously improving decision-making and trading performance.
@@
Self-evolving agents have become a powerful paradigm in medical AI, where adaptability and the ability to evolve are essential for managing the complexity and ever-changing nature of real-world clinical practice. One of the most prominent applications is hospital-scale simulation. For example, Agent Hospital (agent2024hospital, ) creates closed environments with LLM-driven doctors, patients, and nurses, allowing the doctor agent to treat thousands of virtual cases. This process helps these agents autonomously refine and evolve their diagnostic strategies without manual labeling, ultimately achieving strong performance on USMLE-style exams. Similarly, MedAgentSim (MedAgentSim2025, ) integrates an LLM doctor, patient, and tool agent. It records successful consultations as reusable trajectories and employs chain-of-thought reflection and consensus to drive self-evolution, improving success rates over successive interactions. Another example is EvoPatient (du2024evopatient, ) places a doctor agent and a patient agent in continuous dialogue. With each generation, they update their memory with high-quality exchanges: the patient develops more realistic symptom narratives, while the doctor learns to ask sharper questions. Notably, this happens without explicit gradient updates or hand-crafted rewards. Reinforcement learning is also central to building adaptive medical agents. For instance, DoctorAgent-RL (doctoragentRL2025, ) models consultations as a Markov decision process, using a reward function that scores diagnostic accuracy, coverage, and efficiency. This guides policy-gradient updates that help the agent ask more relevant questions and reach correct diagnoses faster than imitation-based approaches, thus achieving self-improvement. In addition, automated architecture-search approaches like Learning to Be a Doctor treat the workflow itself as an evolvable object, iteratively inserting specialist sub-agents or new reasoning hops to cover observed failure modes and improve multimodal diagnostic accuracy (zhuang2025archsearch, ). Finally, beyond clinical decision-making, self-evolving agents have also been extended to biomedical discovery. OriGene (zhang2025origene, ) functions as a virtual disease biologist that evolves by iteratively refining its analytical process. It leverages human and experimental feedback to update core reasoning templates, adjust tool usage strategies, and refine analytical protocols. Similarly, STELLA (jin2025stella, ) is a self-evolving biomedical research agent that improves over time by distilling successful reasoning workflows into reusable templates through its Template Library and expanding its Tool Ocean with external or newly assembled tools to meet emerging analytical needs.
@@
Self-evolving LLM agents have also found strong applications in the education domain. At the learner level, self-evolving agents like the personalized tutor PACE (liu2025pace, ) adjust their prompts based on detailed student profiles and continually refine their questioning during conversations. Meanwhile, an LLM-to-LLM self-play framework generates diverse tutor–student dialogues that further fine-tune the agent, allowing its teaching strategies to evolve both during and after interactions. Another example is MathVC (yue2025mathvc, ), which employs symbolic persona profiles for virtual students and a meta-planner that orchestrates realistic problem-solving stages. This setup enables the agent’s conversational process to evolve step by step toward correct solutions, closely mirroring how collaborative learning naturally unfolds. On the instructor side, self-evolving agent systems like the professional-development platform i-vip (yang2025ivip, ) deploy a team of cooperating LLM agents—a coach, assessor, and feedback generator—that critique and enhance each other’s outputs in real time. These agents adapt their explanations based on teacher-learners’ responses and continue to evolve by incorporating expert feedback after deployment, thereby refining their prompt strategies over time Similarly, EduPlanner (zhang2025eduplanner, ) frames lesson-plan creation as an adversarial loop where a planner’s draft is repeatedly reviewed and refined by evaluator and optimizer agents until it meets diverse educational goals. Similarly, SEFL (zhang2025sefl, ) uses teacher–student self-play to generate large sets of homework–feedback examples, which then fine-tune a lightweight feedback model. This self-evolving process significantly improves the clarity and usefulness of the comments. Collectively, these examples illustrate how self-evolving LLM agents can dynamically adapt to both learners and instructors, driving more personalized, effective, and scalable educational experiences.
@@
Beyond the four major verticals discussed above, self-evolving agents demonstrate broader applicability, delivering superior adaptability and performance in specialized domains where conventional agents often fall short. For instance, Arxiv Copilot (lin2024arxivcopilot, ) learns and adapts by incorporating historical user interactions, including generated answers, research trends, and ideas, into its thought database, enhancing its ability to provide personalized and augmented academic assistance. In a very different context, Voyager (wang2023voyager, ), an agent in the game Minecraft, excels at solving novel tasks from scratch in new worlds through a process of self-evolution. It continually refines its task goals via an automatic curriculum, expands its skill library, and enhances its actions using an iterative prompting mechanism without human intervention. Transitioning to domains that require explicit strategic planning, Agents-of-Change (belle2025agents, ) autonomously refines prompts and rewrites code based on iterative performance analysis and strategic research, thereby helping agents overcome inherent limitations in long-term strategic planning and achieve consistently superior and more coherent gameplay in complex environments like Settlers of Catan. Lastly, in the realm of diplomacy, Richelieu (zhao2024richelieu, ) introduces AI diplomacy agents that can self-evolve through their self-play mechanism, which allows the agent to augment its memory by acquiring diverse experiences without human data, thereby enhancing its strategic planning, reflection, and overall performance in diplomacy activities. While these diverse examples operate in distinct environments—from academic research and virtual game worlds to strategic board games and complex diplomatic negotiations—they all share the fundamental characteristic of leveraging continuous learning, self-refinement, and autonomous adaptation to achieve increasingly sophisticated and effective performance within their respective domains. These diverse examples reinforce the versatility of self-evolving agents, showcasing their growing potential to excel in a wide range of complex, dynamic, and human-like tasks beyond traditional domains.
@@ 7 Evaluation of Self-evolving Agents
Evaluating self-evolving agents presents a unique set of challenges that extend beyond the traditional assessment of static AI systems. Unlike conventional agents typically evaluated on a fixed set of tasks at a single point in time, self-evolving agents are designed to continuously learn, adapt, and improve through ongoing interaction with dynamic environments. Consequently, their evaluation must capture not only immediate task success but also crucial aspects such as adaptation over time, knowledge accumulation and retention, long-term generalization, and the ability to transfer learned skills across sequential or novel tasks, all while mitigating catastrophic forgetting. This demands a fundamental shift from conventional “single-shot” assessments to a longitudinal view of their growth trajectory.
Figure 9: Overview of evaluation angles for self-evolving agents, encompassing core Evaluation Goals and Metrics—such as adaptivity, retention, generalization, safety, and efficiency—and a continuum of Evaluation Paradigms spanning from static assessment to short-term adaptability and long-horizon lifelong learning evaluation.
7.1 Evaluation Goals and Metrics
@@
To effectively evaluate self-evolving agents, we must move beyond traditional metrics and establish a comprehensive framework that captures their dynamic, adaptive, and long-term learning capabilities. A truly capable and desirable self-evolving agent must not only learn and improve but also remember past knowledge, transfer it to new situations, operate sustainably, and behave responsibly. Grounded in these critical requirements for continuous and robust AI, we categorize the key evaluation goals into five core dimensions: Adaptivity, Retention, Generalization, Efficiency, and Safety, as illustrated in Table 5. Each dimension addresses a vital aspect of an agent’s self-evolutionary process, providing a holistic view of its performance.
Adaptivity
@@
Adaptivity serves as a foundational evaluation criterion for any self-evolving agent, measuring its ability to improve performance on in-domain tasks through experience. This dimension focuses on quantifying the learning curve and the extent of performance enhancement as an agent iterates and evolves within a specific domain. Rather than a static success rate, adaptivity is gauged over time, steps, or iterations. Typical metrics include the Success Rate by Iteration Steps hu2024automated ; wang2024agent ; zheng2025lifelongagentbench , which tracks performance in downstream tasks as a function of the agent’s interaction history.
Retention
@@
Retention is a crucial criterion for evaluating the stability of a self-evolving agent’s knowledge base. This dimension specifically focuses on the challenge of catastrophic forgetting, a common issue in lifelong learning where new knowledge acquisition erodes previously learned information, and knowledge retention within extended interactions. Two key metrics can be used to quantify this stability from different perspectives: Forgetting (FGT) and Backward Transfer (BWT) zheng2025lifelong . Specifically, Let $J_{i,t}$ be the performance of LLM agents on task $i$ after completing $t$ tasks. FGT and BWT can be calculated as follows:
$FGT_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}[\displaystyle\max_{j\in\{i,i+1,...,t\}}(\{J_{j,i}\}_{j})-J_{t,i}]\quad BWT_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}(J_{t,i}-J_{i,i})$
@@
FGT evaluates the average accuracy drop on old tasks after an agent learns a new one, thereby measuring whether useful experience is successfully maintained. In contrast, BWT assesses the average accuracy improvement on old tasks due to the experience gained from a new task. A positive BWT indicates that new learning positively benefits old tasks, signifying successful knowledge transfer and a more robust, stable learning process.
Generalization

### v4

@@ 6.2 Specialized Domain Evolution
The power of self-evolving agents extends directly to practical applications like coding, where their ability to autonomously adapt and improve offers a transformative approach to software development. SICA (Robeyns et al., 2025a) demonstrates that a self-improving coding agent can autonomously edit its own codebase and improve its performance on benchmark tasks. EvoMAC (Hu et al., 2024d) introduces a self-evolving paradigm on multi-agent collaboration networks, which automatically optimizes individual agent prompts and multi-agent workflows, significantly improving code generation performance by overcoming the limitations of manually designed systems. AgentCoder (Huang et al., 2024) also focuses on a multi-agent code generation framework that self-evolves through iterative refinement. A programmer agent continuously improves code based on feedback from a test executor agent, validated against independent test cases from a test designer, significantly boosting effectiveness and efficiency. Zhang et al. (Zhang et al., 2025b) enable LLM agents to continuously evolve by filtering high-quality answers, stratifying earned experiences by difficulty, and adaptively selecting demonstrations from self-generated data, leading to significant performance improvements and the construction of ML libraries. While these instances differ in their specific mechanisms—ranging from single-agent self-editing to complex multi-agent collaborative networks and experience-based learning—they commonly share the core principle of iterative self-improvement and autonomous adaptation to enhance coding capabilities. These advancements highlight how self-evolving agents can dramatically enhance coding efficiency and code quality by continuously learning and optimizing.
@@
Self‑evolving GUI agents extend LLM capabilities from pure text reasoning to direct manipulation of desktop, web, and mobile interfaces, where they must cope with large discrete action spaces, heterogeneous layouts, and partial visual observability. Yuan et al. couple pixel‑level vision with self‑reinforcement, enabling the agent to iteratively refine click–type grounding accuracy without additional human labels (Yuan et al., 2025c). On real desktop software, the Navi agent from WindowsAgentArena replays and critiques its own failure trajectories, ultimately doubling its task‑completion rate across 150 Windows challenges (Bonatti et al., 2024). For open‑web automation, WebVoyager fuses screenshot features with chain‑of‑thought reflection; successive self‑fine‑tuning raises its end‑to‑end success on unseen sites from 30 % to 59 % (He et al., 2024), while ReAP adds episodic memories of past outcomes, recovering a further 29‑percentage‑point margin on previously failed queries (Azam et al., 2025). Beyond RL and memory, AutoGUI continuously mines functionality annotations from live interfaces to expand a reusable skill library each training cycle (Li et al., 2025a), and MobileUse deploys a hierarchical self‑reflection stack that monitors, verifies, and revises smartphone actions in situ (Li et al., 2025c). Collectively, these systems epitomize the full triad of self‑evolution— what evolves (grounding modules, skill memories), when it evolves (offline consolidation vs. online reflection), and how it evolves (reinforcement learning, synthetic data, hierarchical monitoring)—charting a path toward universally competent interface agents.
@@
The primary bottleneck in customizing agents for specialized domains like financial tasks lies in efficiently constructing and integrating a domain-specific knowledge base into the agent’s learning process—a challenge that can be effectively mitigated by incorporating self-evolving mechanisms. QuantAgent (Wang et al., 2024d) proposed a two-layer framework that iteratively refines the agent’s responses and automatically enhances its domain-specific knowledge base using feedback from simulated and real-world environments. This iterative process helps the agent progressively approximate optimal behavior, reduces reliance on costly human-curated datasets, and demonstrably improves its predictive accuracy and signal quality in trading tasks. TradingAgents (Xiao et al., 2024) incorporates dynamic processes such as reflection, reinforcement learning, and a feedback loop from real-world trading results, alongside collaborative debates, to continuously refine its strategies and enhance trading performance. These developments underscore the potential of self-evolving agents to revolutionize the financial domain by autonomously building domain expertise, adapting to dynamic market conditions, and continuously improving decision-making and trading performance.
@@
Self-evolving agents have become a powerful paradigm in medical AI, where adaptability and the ability to evolve are essential for managing the complexity and ever-changing nature of real-world clinical practice. One of the most prominent applications is hospital-scale simulation. For example, Agent Hospital (Li et al., 2024a) creates closed environments with LLM-driven doctors, patients, and nurses, allowing the doctor agent to treat thousands of virtual cases. This process helps these agents autonomously refine and evolve their diagnostic strategies without manual labeling, ultimately achieving strong performance on USMLE-style exams. Similarly, MedAgentSim (Almansoori et al., 2025b) integrates an LLM doctor, patient, and tool agent. It records successful consultations as reusable trajectories and employs chain-of-thought reflection and consensus to drive self-evolution, improving success rates over successive interactions. Another example is EvoPatient (Du et al., 2024) places a doctor agent and a patient agent in continuous dialogue. With each generation, they update their memory with high-quality exchanges: the patient develops more realistic symptom narratives, while the doctor learns to ask sharper questions. Notably, this happens without explicit gradient updates or hand-crafted rewards. Reinforcement learning is also central to building adaptive medical agents. For instance, DoctorAgent-RL (Feng et al., 2025c) models consultations as a Markov decision process, using a reward function that scores diagnostic accuracy, coverage, and efficiency. This guides policy-gradient updates that help the agent ask more relevant questions and reach correct diagnoses faster than imitation-based approaches, thus achieving self-improvement. In addition, automated architecture‑search approaches like Learning to Be a Doctor treat the workflow itself as an evolvable object, iteratively inserting specialist sub‑agents or new reasoning hops to cover observed failure modes and improve multimodal diagnostic accuracy (Zhuang et al., 2025). Finally, beyond clinical decision-making, self-evolving agents have also been extended to biomedical discovery. OriGene (Zhang et al., 2025t) functions as a virtual disease biologist that evolves by iteratively refining its analytical process. It leverages human and experimental feedback to update core reasoning templates, adjust tool usage strategies, and refine analytical protocols. Similarly, STELLA (Jin et al., 2025) is a self-evolving biomedical research agent that improves over time by distilling successful reasoning workflows into reusable templates through its Template Library and expanding its Tool Ocean with external or newly assembled tools to meet emerging analytical needs.
@@
Self-evolving LLM agents have also found strong applications in the education domain. At the learner level, self-evolving agents like the personalized tutor PACE (Liu et al., 2025c) adjust their prompts based on detailed student profiles and continually refine their questioning during conversations. Meanwhile, an LLM-to-LLM self-play framework generates diverse tutor–student dialogues that further fine-tune the agent, allowing its teaching strategies to evolve both during and after interactions. Another example is MathVC (Yue et al., 2025), which employs symbolic persona profiles for virtual students and a meta-planner that orchestrates realistic problem-solving stages. This setup enables the agent’s conversational process to evolve step by step toward correct solutions, closely mirroring how collaborative learning naturally unfolds. On the instructor side, self-evolving agent systems like the professional-development platform i‑vip (Yang et al., 2025b) deploy a team of cooperating LLM agents—a coach, assessor, and feedback generator—that critique and enhance each other’s outputs in real time. These agents adapt their explanations based on teacher-learners’ responses and continue to evolve by incorporating expert feedback after deployment, thereby refining their prompt strategies over time Similarly, EduPlanner (Zhang et al., 2025p) frames lesson-plan creation as an adversarial loop where a planner’s draft is repeatedly reviewed and refined by evaluator and optimizer agents until it meets diverse educational goals. Similarly, SEFL (Zhang et al., 2025k) uses teacher–student self-play to generate large sets of homework–feedback examples, which then fine-tune a lightweight feedback model. This self-evolving process significantly improves the clarity and usefulness of the comments. Collectively, these examples illustrate how self-evolving LLM agents can dynamically adapt to both learners and instructors, driving more personalized, effective, and scalable educational experiences.
@@
Beyond the four major verticals discussed above, self-evolving agents demonstrate broader applicability, delivering superior adaptability and performance in specialized domains where conventional agents often fall short. For instance, Arxiv Copilot (Lin et al., 2024) learns and adapts by incorporating historical user interactions, including generated answers, research trends, and ideas, into its thought database, enhancing its ability to provide personalized and augmented academic assistance. In a very different context, Voyager (Wang et al., 2023a), an agent in the game Minecraft, excels at solving novel tasks from scratch in new worlds through a process of self-evolution. It continually refines its task goals via an automatic curriculum, expands its skill library, and enhances its actions using an iterative prompting mechanism without human intervention. Transitioning to domains that require explicit strategic planning, Agents-of-Change (Belle et al., 2025) autonomously refines prompts and rewrites code based on iterative performance analysis and strategic research, thereby helping agents overcome inherent limitations in long-term strategic planning and achieve consistently superior and more coherent gameplay in complex environments like Settlers of Catan. Lastly, in the realm of diplomacy, Richelieu (Zhao et al., 2024d) introduces AI diplomacy agents that can self-evolve through their self-play mechanism, which allows the agent to augment its memory by acquiring diverse experiences without human data, thereby enhancing its strategic planning, reflection, and overall performance in diplomacy activities. While these diverse examples operate in distinct environments—from academic research and virtual game worlds to strategic board games and complex diplomatic negotiations—they all share the fundamental characteristic of leveraging continuous learning, self-refinement, and autonomous adaptation to achieve increasingly sophisticated and effective performance within their respective domains. These diverse examples reinforce the versatility of self-evolving agents, showcasing their growing potential to excel in a wide range of complex, dynamic, and human-like tasks beyond traditional domains.
@@ 7 Evaluation of Self-evolving Agents
Evaluating self-evolving agents presents a unique set of challenges that extend beyond the traditional assessment of static AI systems. Unlike conventional agents typically evaluated on a fixed set of tasks at a single point in time, self-evolving agents are designed to continuously learn, adapt, and improve through ongoing interaction with dynamic environments. Consequently, their evaluation must capture not only immediate task success but also crucial aspects such as adaptation over time, knowledge accumulation and retention, long-term generalization, and the ability to transfer learned skills across sequential or novel tasks, all while mitigating catastrophic forgetting. This requires a fundamental shift from conventional “single-shot” scoring to a longitudinal, cost-aware trajectory view.
Figure 9: An overview of evaluation goals and paradigms for self-evolving agents. Left: Evaluation Goals and Metrics. We summarize five core objectives guiding agent evaluation: adaptivity, retention, generalization, efficiency, and safety, which reflect how agents should evolve and perform over time. Right: Evaluation Paradigm. Evaluation methods are organized along an increasing temporal scale: from static assessment of fixed capabilities (e.g., external task-solving and component-level evaluation), to short-horizon adaptive assessment capturing within-task adaptation (e.g., augmented or dynamic benchmarks), and finally to long-horizon lifelong learning assessment that examines continual improvement under evolving or lifelong benchmarks. Together, these axes link the goals of self-evolution with the evaluation settings used to measure them, forming a unified framework for assessing agent adaptivity and sustainability. See Table 6 for metric definitions, Table 10 for protocol summary.
7.1 Evaluation Goals, Metrics, and Benchmark Coverage
@@
To effectively evaluate self-evolving agents, we must move beyond traditional metrics and establish a comprehensive framework that captures their dynamic, adaptive, and long-term learning capabilities. A truly capable and desirable self-evolving agent must not only learn and improve but also remember past knowledge, transfer it to new situations, operate sustainably, and behave responsibly. Grounded in these critical requirements for continuous and robust AI, we categorize the key evaluation goals into five core dimensions: Adaptivity, Retention, Generalization, Efficiency, and Safety, as illustrated in Table 6. Each dimension addresses a vital aspect of an agent’s self-evolutionary process and is assessed through its corresponding metrics and benchmark coverage analysis. While Table 7 provides a comprehensive catalog of evaluation resources, we synthesize these limitations to map coverage gaps across these five-goal framework and identify directions for trajectory-centric, cost-aware assessment at Table  8.
7.1.1 Adaptivity
Goals & Metrics
@@ 7.1.1 Adaptivity
Adaptivity serves as a foundational evaluation criterion for any self-evolving agent, measuring its ability to improve performance on in-domain tasks through experience. This dimension focuses on quantifying the learning curve and the extent of performance enhancement as an agent iterates and evolves within a specific domain. Rather than a static success rate, adaptivity is gauged over time, steps, or iterations. Typical metrics include the Success Rate by Iteration Steps (Hu et al., 2024c; Wang et al., 2024j; Zheng et al., 2025b), which tracks performance in downstream tasks as a function of the agent’s interaction history.
Coverage Gaps
Although adaptivity has the richest benchmark ecosystem—spanning code generation (SWE-bench (Jimenez et al., 2023), MLE-Bench (Chan et al., 2024)), web navigation (WebArena (Zhou et al., 2023), WebShop (Yao et al., 2022)), and general reasoning (GAIA (Mialon et al., 2023), AgentBench (Liu et al., 2023b) evaluations remain constrained by practical design choices. ScienceAgentBench limits tasks to a single programming language and imposes execution-time bounds, excluding computationally intensive scientific workflows (Chen et al., 2024e). MLE-Bench offers well-structured tasks but diverges from authentic research scenarios where problem formulation is itself part of the challenge (Chan et al., 2024). These simplifications permit controlled in-domain assessment but confine evaluation to narrow, static task distributions. Moreover, most benchmarks measure improvement under pre-specified learning protocols (e.g., fixed iteration budgets or predetermined replay schedules) rather than assessing whether agents autonomously discover effective adaptation strategies. AgentBench further reports weaknesses in sustained reasoning and strategic decision-making (Liu et al., 2023b), yet evaluation horizons remain limited to short-term improvement curves.
7.1.2 Retention
Goals & Metrics
@@ 7.1.2 Retention
Retention is a crucial criterion for evaluating the stability of a self-evolving agent’s knowledge base. This dimension specifically focuses on the challenge of catastrophic forgetting, a common issue in lifelong learning where new knowledge acquisition erodes previously learned information, and knowledge retention within extended interactions. Two key metrics can be used to quantify this stability from different perspectives: Forgetting (FGT) and Backward Transfer (BWT) (Zheng et al., 2025c). Specifically, let $J_{i,t}$ denote the performance on task $i$ after completing $t$ tasks:
$\mathrm{FGT}_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}\!\Big[\max_{j\in\{i,\ldots,t\}}J_{i,j}-J_{i,t}\Big],\quad\mathrm{BWT}_{t}=\frac{1}{t-1}\sum_{i=1}^{t-1}\big(J_{i,t}-J_{i,i}\big).$
@@
A positive BWT indicates that new learning positively benefits old tasks, signifying successful knowledge transfer and a more robust, stable learning process.
Coverage Gaps
Retention remains the most underserved dimension. Current memory mechanisms struggle significantly with dynamic state updates and maintaining consistency across extended interactions (Hu et al., 2025). Experience replay exhibits an inherent tension: while replay buffers can improve learning, scaling them beyond optimal thresholds triggers performance degradation through context overflow and resource exhaustion (Zheng et al., 2025b). Economic and computational constraints limit evaluation robustness, with some benchmarks conducting minimal repetitions that may not capture stochastic variation (Castillo-Bolado et al., 2024). Most critically, the overwhelming majority of existing benchmarks adopt episodic evaluation where agent state resets between tasks, fundamentally precluding measurement of knowledge accumulation or degradation—precisely the phenomena that distinguish self-evolving agents from static systems.
7.1.3 Generalization
Goals & Metrics